How many cubic centimeters of water will fill this glass
the height is 20 cm and the radius is 3 cm

Answer by jim_thompson5910(35256)

V = pi*r^2*h

V = pi*(3)^2*(20)

V = pi*9*20

V = pi*180

V = 180pi <--- There's the exact volume

V = 180(3.14159)

V = 565.4862 <--- There's the approximate volume